# Short Film Dataset (SFD): A Benchmark for Story-Level Video Understanding

### v1

@@ Appendix I Prompts
We provide the prompts for the data generation process: question generation prompt (see Figure 8), distractor generation prompt (see Figure 9), and question classification prompt (see Figure 10). As detailed in Section 2.2 , those prompts were used with GPT-4 (’gpt-4-turbo-2024-04-09’) through the OpenAI API.
![[Uncaptioned image]](x20.png)
![[Uncaptioned image]](x21.png)
![[Uncaptioned image]](x22.png)
@@
This dataset and benchmark are created to assess long-form video understanding capabilities of modern multimodal systems. It is different from other video question-answering datasets as it features longer videos and associated story-related questions. Additionally, it stands out as a dataset that is available online and not exposed to current large language models (limited data leakage issues).
@@
This dataset was created by Ridouane Ghermi††footnotemark: , Xi Wang††footnotemark: , Vicky Kalogeiton††footnotemark: , and Ivan Laptev999MBZUAI.
@@
Our dataset instance consists of a short movie (around 5-20 mins), some metadata (e.g. movie title, logline, synopsis, country of origin, language, release year), and a list of multiple-choice questions and answers.
@@
There are 1,078 movies and 4,885 questions. Each question comprises one correct answer and four wrong answers.
@@
The dataset is not a sample of instances from a larger set.
@@
Each instance includes a movie as a video file. Metadata comes as .csv files, in the form of text. We also provide processed features.
@@
We provide a dataframe that links question ids and video ids.
@@
SFD is a test set only.
@@
All movies are available on YouTube; We only provide the corresponding URLs. It should remain constant over time as it comes from YouTube channels that curate a movie catalogue. The dataset is distributed under the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license.
@@
Movies are available online and downloaded from YouTube using yt-dlp, along with metadata. Multiple-choice questions are generated by a commercial LLM, GPT-4, and further manually curated.
@@
Video URLs are scraped using the official YouTube API. Videos and metadata are downloaded using yt-dlp. Questions are generated using GPT-4 through the official OpenAI API and further refined by manual curation.
@@
We extracted several annotations from the dataset, see Sections C.
@@
The dataset is accompanied by two tasks: multiple-choice and open-ended question answering.
@@
Yes, we will make this available on the dataset GitHub repository.
@@
The dataset is distributed on the HuggingFace platform and as a GitHub repository. We provide video URLs and annotations. Additionally, we provide code to download videos, pre-process the data and run baselines.
@@
The dataset is a test benchmark, it will remain stable to ensure fair evaluation of future methods.
@@
If others want to contribute to the dataset, by adding more tasks for example, we would gladly host them in our repository. Also, we will provide as many details and tools as possible to reproduce our annotation process and create more tasks.

### v2

@@ Appendix H Prompts
We provide the prompts for the data generation process: question generation prompt (see Table 14), distractor generation prompt (see Table 15), and question classification prompt (see Table 16). As detailed in Section 3.2, those prompts were used with GPT-4 (’gpt-4-turbo-2024-04-09’) through the OpenAI API.
![[Uncaptioned image]](x13.png)
![[Uncaptioned image]](x14.png)
![[Uncaptioned image]](x15.png)
Appendix I Samples from SF20K-Test
In the following pages, we show some samples from SF20K-Test. For each sample, we display the title, the logline, the synopsis, a few frames, some subtitles, and the designed multiple-choice questions. See Figures 10, 11,12, 13.
@@ Appendix I Samples from SF20K-Test
This dataset is created to improve long-form video understanding capabilities of modern multimodal systems. It comes with a benchmark to assess current performance. This video question-answering dataset differs from previous ones as it features longer videos and associated story-related questions. Additionally, it stands out as a dataset that is available online and not exposed to current large language models (which limits data contamination issues).
@@
This dataset was created by Ridouane Ghermi††footnotemark: , Xi Wang††footnotemark: , Vicky Kalogeiton††footnotemark: , and Ivan Laptev121212MBZUAI.
@@
A single instance consists of a short movie (around 5-50 mins), and corresponding metadata (i.e. movie title, logline, country of origin, language, release year).
@@
There are 20,143 movies.
@@
The dataset is new, this is not a sample from a larger set.
@@
Each instance consists of a movie (video file), metadata (csv file), and extracted annotations (i.e. shot boundaries, subtitles, captions, and face tracks)
@@
Not applicable.
@@
SF20K contains a train set and a test set, called SF20K-Test. Subsequently, SF20K-Test contains two subsets: SF20K-Test-Silent and SF20K-Test-Expert.
@@
All movies are available on YouTube and Vimeo. We only provide URLs. It should remain constant over time as it comes from YouTube and Vimeo channels that curate a movie catalogue. The dataset is distributed under the CC BY-NC-SA 4.0 (Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International) license.
@@
Movies and metadata are available online, and downloaded directly from YouTube and Vimeo.
@@
Video URLs are scraped using the official APIs. Videos and metadata are downloaded using yt-dlp.
@@
We extracted several annotations from the dataset, see Sections B.
@@
The dataset is used to train video question answering models.
@@
Yes.
@@
The dataset is distributed on the HuggingFace platform and as a GitHub repository. We provide video URLs and annotations. Additionally, we provide code to download videos, pre-process the data, run baselines, and fine-tune models.
@@
The dataset might be updated to increase its size and diversity. However, the test set will remain stable to ensure fair evaluation of future methods.
@@
If others want to contribute to the dataset, by adding videos or labels for example, we would gladly host them in our repository. Also, we will provide as many details and tools as possible to reproduce our annotation process and create more tasks.